# Changzhou, 213164, China

### v1

@@ 1 Introduction
Searches for the triplet scalar have been extensively conducted at the LHC, as its charged components, especially the doubly charged ones are predicted not only in the type-II seesaw, but also in a variety of other beyond SM models such as the left-right symmetric models Pati:1974yy ; Mohapatra:1974hk ; Senjanovic:1975rk , Zee-Babu model Zee:1985id ; Babu:1988ki , Georgi-Machacek model Chanowitz:1985ug ; Georgi:1985nv , etc. Pair production and associated production of these exotic scalars from Drell-Yan processes are commonly considered, and multileptonic final states are searched for as the most promising signals. The obtained experimental bounds for the doubly charged Higgs bosons are generally sensitive to different parametric regions of the model, mainly concerned with the triplet scalar VEV $v_{\Delta}$ and the mass spectrum of the charged and neutral components. It is the value of $v_{\Delta}$ that largely determines decay modes of the doubly charged scalars. Assuming $v_{\Delta}$ far less than $10^{-4}$ GeV which leads to dominant dileptonic decay channels $\Delta^{\pm\pm}\to\ell^{\pm}\ell^{{}^{\prime}\pm}$, ATLAS collaboration derived a stringent lower limit for the doubly charged scalar mass at 95% C.L. $m_{\Delta}>1080$ GeV ATLAS:2022pbd . With a larger $v_{\Delta}>10^{-4}$ GeV, the dileptonic decay channels are highly suppressed due to the increasing $v_{\Delta}$ and only diboson channels $\Delta^{\pm\pm}\to W^{\pm}W^{\pm}$ are relevant, relatively weaker bounds are obtained at $200\sim 220$ GeV for a degenerate spectrum between the singly and doubly charged Higgs ATLAS:2018ceg . For a non-degenerate spectrum with a mass difference of up to 100 GeV, lower limits of 350 and 230 GeV are obtained for pair and associated production, respectively ATLAS:2021jol . Although most current experimental searches assumed degeneracy or a tiny mass splitting, it is allowed to be as large as $\Delta m\sim 40$ GeV for hundreds-GeV triplet Higgs if considering the electroweak precision data (EWPD) due to the fact that heavy scalars contribution to the oblique parameter $T$ is compensated by the mass difference among the triplet components Melfo:2011nx ; Chun:2012jw . In addition, for a moderate mass separation $\Delta m\gtrsim 5$ GeV or even larger splitting, cascade decay channels $\Delta^{\pm\pm}\to\Delta^{\pm}W^{\pm*}\to\Delta^{0}W^{\pm*}W^{\pm*}$ become dominant over the dileptonic and dibosonic ones if $v_{\Delta}$ is not very far from $10^{-4}$ GeV by orders (either larger or smaller) Melfo:2011nx . Hence, when the decay of the new scalars is dominated by different channels, quite distinct signatures appear in collider phenomenology. And due to the large hadronic background at the LHC, a tiny $v_{\Delta}$ is typically assumed for a dominant dileptonic mode in experimental searches as well as phenomenological studies.
@@ 4 Search strategies and results
To arrive at an effective search strategy for the fully leptonic channel from the doubly charged scalar at the LHC, we perform simulations using MadGraph5_aMC@NLO (v3.5.6) Alwall2014 for parton-level events generation and calculations for cross sections. Parton showering and detector simulation are realized by Pythia-8.2 Sjostrand:2014zea and Delphes-3.5.0 deFavereau:2013fsa embedded in CheckMATE2 Dercks2017 for further analysis and event selection. For model files, we use the Typeiiseesaw Universal Feynrules Output (UFO) libraries that are developed in Fuks:2019clu .
@@
Figure 5: Integrated luminosities required to reach $2\sigma$ exclusion searching for fully elastic photon fusion $pp\to p(\gamma\gamma\to\Delta^{++}\Delta^{--}\to\ell^{+}\ell^{+}\ell^{-}\ell^{-})p$ at the 14 TeV LHC. Systematic uncertainty $\beta$ is assumed as 5%.
@@
Figure 6: Branching ratios that can be excluded at 95% C.L. via searching for fully elastic photon fusion $pp\to p(\gamma\gamma\to\Delta^{++}\Delta^{--}\to\ell^{+}\ell^{+}\ell^{-}\ell^{-})p$ at the 14 TeV LHC, shown in black ($\mathcal{L}=36.1\text{fb}^{-1}$), blue ($\mathcal{L}=100\text{fb}^{-1}$) and red ($\mathcal{L}=3\text{ab}^{-1}$) solid lines. Systematic uncertainty $\beta$ is assumed as 5%. Figure (a) and (b) are respectively for branching ratios of $\Delta^{\pm\pm}\to e^{\pm}e^{\pm}$ and $\mu^{\pm}\mu^{\pm}$. Dashed lines are bounds from the ATLAS experiment ATLAS:2012hi ; ATLAS:2014kca ; ATLAS:2017xqs .
@@ 5 Conclusion
We study the sensitivities for pair production of doubly charged scalars within the type-II seesaw mechanism (with a degenerate mass spectrum of the scalar components) through EPA photon fusion at the 14 TeV LHC: $pp\to p(\gamma\gamma\to\Delta^{++}\Delta^{--})p$. Two kinds of neutrino mass spectra, normal and inverted hierarchies, are considered. In the case of normal hierarchy, final states of $\mu^{+}\mu^{+}\mu^{-}\mu^{-}$ are studied as the signal process with integrated luminosities of 36.1 fb-1, 100 fb-1 and 3 ab-1, the $2\sigma$ exclusion limits on $m_{\Delta}$ can reach around 430, 520 and 1000 GeV respectively, for branching ratio of $\Delta^{\pm\pm}\to\mu^{\pm}\mu^{\pm}\sim$ 25%, which is set by adopting the best-fit values of PMNS matrix under NH spectrum. And for the case of inverted hierarchy and a largely fixed branching ratio $\Delta^{\pm\pm}\to e^{\pm}e^{\pm}\sim$ 47%, final states of $e^{+}e^{+}e^{-}e^{-}$ are studied as signal under the same luminosities with the exclusion bound on $m_{\Delta}$ reaching about 730, 880 and 1300 GeV, corresponding to the above luminosities.
@@ 6 Acknowledgments
This work is supported by the National Natural Science Foundation of China under Grant No. 12405118 and the Natural Science Foundation of Jiangsu Province under Grant No. BK20230623.

### v2

@@ 1 Introduction
Searches for the triplet scalar at colliders and studies on the relevant features have been extensively conducted Ghosh:2017pxl ; Crivellin:2018ahj ; BhupalDev:2018tox ; Altakach:2022hgn ; Das:2023tna ; Das:2024kyk ; Arroyo-Urena:2025boh , as its charged components, especially the doubly charged ones are predicted not only in the type-II seesaw, but also in a variety of other beyond SM models such as the left-right symmetric models Pati:1974yy ; Mohapatra:1974hk ; Senjanovic:1975rk , Zee-Babu model Zee:1985id ; Babu:1988ki , Georgi-Machacek model Chanowitz:1985ug ; Georgi:1985nv , etc. Pair production and associated production of these exotic scalars from Drell-Yan processes are commonly considered, and multileptonic final states are searched for as the most promising signals. The obtained experimental bounds for the doubly charged Higgs bosons are generally sensitive to different parametric regions of the model, mainly concerned with the triplet scalar VEV $v_{\Delta}$ and the mass spectrum of the charged and neutral components. It is the value of $v_{\Delta}$ that largely determines decay modes of the doubly charged scalars. Assuming $v_{\Delta}$ far less than $10^{-4}$ GeV which leads to dominant dileptonic decay channels $\Delta^{\pm\pm}\to\ell^{\pm}\ell^{{}^{\prime}\pm}$, ATLAS collaboration derived a stringent lower limit for the doubly charged scalar mass at 95% C.L. $m_{\Delta}>1080$ GeV ATLAS:2022pbd . With a larger $v_{\Delta}>10^{-4}$ GeV, the dileptonic decay channels are highly suppressed due to the increasing $v_{\Delta}$ and only diboson channels $\Delta^{\pm\pm}\to W^{\pm}W^{\pm}$ are relevant, relatively weaker bounds are obtained at $200\sim 220$ GeV for a degenerate spectrum between the singly and doubly charged Higgs ATLAS:2018ceg . For a non-degenerate spectrum with a mass difference of up to 100 GeV, lower limits of 350 and 230 GeV are obtained for pair and associated production, respectively ATLAS:2021jol . Although most current experimental searches assumed degeneracy or a tiny mass splitting, it is allowed to be as large as $\Delta m\sim 40$ GeV for hundreds-GeV triplet Higgs if considering the electroweak precision data (EWPD) due to the fact that heavy scalars contribution to the oblique parameter $T$ is compensated by the mass difference among the triplet components Melfo:2011nx ; Chun:2012jw . In addition, for a moderate mass separation $\Delta m\gtrsim 5$ GeV or even larger splitting, cascade decay channels $\Delta^{\pm\pm}\to\Delta^{\pm}W^{\pm*}\to\Delta^{0}W^{\pm*}W^{\pm*}$ become dominant over the dileptonic and dibosonic ones if $v_{\Delta}$ is not very far from $10^{-4}$ GeV by orders (either larger or smaller) Melfo:2011nx . Hence, when the decay of the new scalars is dominated by different channels, quite distinct signatures appear in collider phenomenology. And due to the large hadronic background at the LHC, a tiny $v_{\Delta}$ is typically assumed for a dominant dileptonic mode in experimental searches as well as phenomenological studies.
@@ 4 Search strategies and results
To arrive at an effective search strategy for the fully leptonic channel from the doubly charged scalar at the LHC, we perform simulations using MadGraph5_aMC@NLO (v3.5.6) Alwall2014 for parton-level events generation and calculations for cross sections. Parton showering and detector simulation are realized by Pythia-8.2 Sjostrand:2014zea and Delphes-3.5.0 deFavereau:2013fsa embedded in CheckMATE2 Dercks2017 for further analysis and event selection. For model files, we use the Typeiiseesaw Universal Feynrules Output (UFO) libraries that are developed in Fuks:2019clu . MMHT+LUXqed NLO PDF sets Buckley:2014ana ; Manohar:2016nzj ; Manohar:2017eqh ; Harland-Lang:2019pla are used for event generation of photon fusion. More discussion about $\gamma$-PDFs can be found below in the end of this section when we present our results and compare them with ones using other $\gamma$-PDF.
@@
Figure 5: Integrated luminosities required to reach $2\sigma$ exclusion searching for fully elastic photon fusion $pp\to p(\gamma\gamma\to\Delta^{++}\Delta^{--}\to\ell^{+}\ell^{+}\ell^{-}\ell^{-})p$ at the 14 TeV LHC. Systematic uncertainty $\beta$ is assumed as 5%. Solid lines correspond to results using MMHT+LUXqed NLO PDF sets, while dotdashed lines correspond to results considering corrections using $\gamma$-UPC package Shao:2022cly as PDFs for photon fusion.
@@
Figure 6: Branching ratios that can be excluded at 95% C.L. via searching for fully elastic photon fusion $pp\to p(\gamma\gamma\to\Delta^{++}\Delta^{--}\to\ell^{+}\ell^{+}\ell^{-}\ell^{-})p$ at the 14 TeV LHC, shown in black ($\mathcal{L}=36.1\text{fb}^{-1}$), blue ($\mathcal{L}=100\text{fb}^{-1}$) and red ($\mathcal{L}=3\text{ab}^{-1}$) solid lines. Systematic uncertainty $\beta$ is assumed as 5%. Figure (a) and (b) are respectively for branching ratios of $\Delta^{\pm\pm}\to e^{\pm}e^{\pm}$ and $\mu^{\pm}\mu^{\pm}$. Solid lines correspond to results using MMHT+LUXqed NLO PDF sets, while dotdashed lines correspond to results considering corrections using $\gamma$-UPC package Shao:2022cly as PDFs for photon fusion. Dashed lines are bounds from the ATLAS experiment ATLAS:2012hi ; ATLAS:2014kca ; ATLAS:2017xqs .
@@
It should be noted that development on $\gamma$-PDF and its implementation in automated event generation will improve the accuracy of phenomenological studies related to photon fusion at hadron colliders, one of which is the recently developed $\gamma$-UPC package Shao:2022cly which incorporates effects including hadronic survival probabilities. As a conservative estimate, if we make a correction for the photon fusion with $\gamma$-UPC by adopting 75% of the cross sections given by MMHT+LUXqed NLO PDF sets, the exclusion mass bounds related to branching ratios are loosened but can still be comparable with current experimental limits. These corrected contours are shown in Figure. 5 and 6 as dotdashed lines with the same colors corresponding to each of the above luminosities.
@@ 5 Conclusion
We study the sensitivities for pair production of doubly charged scalars within the type-II seesaw mechanism (with a degenerate mass spectrum of the scalar components) through EPA photon fusion at the 14 TeV LHC: $pp\to p(\gamma\gamma\to\Delta^{++}\Delta^{--})p$. Two kinds of neutrino mass spectra, normal and inverted hierarchies, are considered. In the case of normal hierarchy, final states of $\mu^{+}\mu^{+}\mu^{-}\mu^{-}$ are studied as the signal process with integrated luminosities of 36.1 fb-1, 100 fb-1 and 3 ab-1, the $2\sigma$ exclusion limits on $m_{\Delta}$ can reach around 430, 520 and 1000 GeV respectively, for branching ratio of $\Delta^{\pm\pm}\to\mu^{\pm}\mu^{\pm}\sim$ 25%, which is set by adopting the best-fit values of PMNS matrix under NH spectrum. And for the case of inverted hierarchy and a largely fixed branching ratio $\Delta^{\pm\pm}\to e^{\pm}e^{\pm}\sim$ 47%, final states of $e^{+}e^{+}e^{-}e^{-}$ are studied as signal under the same luminosities with the exclusion bounds on $m_{\Delta}$ reaching about 730, 880 and 1300 GeV, corresponding to the above luminosities. The results for exclusion bounds can also be comparable to experimental limits if we take into consideration of the correction by using $\gamma$-UPC package as PDFs for photon fusion.
@@ 6 Acknowledgments
We are grateful to David d’Enterria and Hua-Sheng Shao for their helpful discussions on $\gamma$-PDFs. This work is supported by the National Natural Science Foundation of China under Grant No. 12405118 and the Natural Science Foundation of Jiangsu Province under Grant No. BK20230623.